OpenStudy (anonymous):

cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (anonymous):

Is this equation or identity?

OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation.

ganeshie8 (ganeshie8):

\(\cot x \sec^4 x = \cot x + 2 \tan x + \tan^3 x \)

ganeshie8 (ganeshie8):

its like above ?

OpenStudy (anonymous):

huh?

ganeshie8 (ganeshie8):

*the equation u seeing in ur assessment sheet, is it like the one ive posted above ?

OpenStudy (anonymous):

its tan cubed x at the end other then that yes

ganeshie8 (ganeshie8):

ok thnks :) pick the side that has more terms, and do SOMETHING and try getting to the other side

ganeshie8 (ganeshie8):

here, the right side has more terms, right ?

OpenStudy (anonymous):

what? the irght side has more terms yeah

ganeshie8 (ganeshie8):

ya so we start with that side, and work, and prove that it equals left side.

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

\(\cot x + 2\tan x + \tan^3 x\) \(\frac{1}{\tan x} + 2\tan x + \tan^3 x\) \(\frac{1+ 2\tan^2 x + \tan^4 x}{\tan x}\) \(\frac{1+ \tan^2 x + \tan^2x + \tan^4 x}{\tan x}\) \(\frac{\sec^2 x + \tan^2x + \tan^4 x}{\tan x}\) \(\frac{\sec^2 x + \tan^2x(1 + \tan^2 x)}{\tan x}\) \(\frac{\sec^2 x + \tan^2x(\sec^2 x)}{\tan x}\) \(\frac{\sec^2 x(1 + \tan^2x)}{\tan x}\) \(\frac{\sec^2 x(\sec^2 x)}{\tan x}\) \(\frac{\sec^4 x}{\tan x}\) \(\cot x\sec^4 x\) = LEFT HAND SIDE

ganeshie8 (ganeshie8):

thats the complete solution; see if it makes sense

OpenStudy (anonymous):

good solve ganeshie.,...........

OpenStudy (anonymous):

How did you get rid of the tan^4x? @ganeshie8?

OpenStudy (mayaal):

@ganeshie8 i dont understand ur solution

ganeshie8 (ganeshie8):

hmm which line ?

OpenStudy (mayaal):

the 3rd line from the end.

ganeshie8 (ganeshie8):

\[\large \frac{\sec^2 x + \tan^2x(\sec^2 x)}{\tan x}\]

ganeshie8 (ganeshie8):

you're fine, till this line ?

OpenStudy (mayaal):

yes

ganeshie8 (ganeshie8):

good, next factor out `sec^2x` from both terms, you get : \[\large \frac{\sec^2 x(1 + \tan^2x)}{\tan x}\]

ganeshie8 (ganeshie8):

still fine ?

OpenStudy (mayaal):

oh,ok.so u factored out the sec^2x from the whole numerator?

ganeshie8 (ganeshie8):

exactly !

OpenStudy (mayaal):

great!thnku very much @ganeshie8

ganeshie8 (ganeshie8):

np :)